Calculate contribution margin ratio and required sales in dollars to breakeven

Eastland Corp had total variable costs of $167,706,total fixed costs of $118,500 and total revenues of $254,100


Calculate contribution margin ratio    


Compute the required sales in dollars to breakeven (round answer to 0 decimal places)
Answer #1

Calculation of Contribution margin ratio = Contribution margin / Sales 

                                                               = 86,394 / 2,54,100

                                                               = 34%


Contributin margin = total revenue - total variable cost

                              =  2,54,100 - 1,67,706

                              =  86,394


Breakeven sales = Fixed Cost / Contribution margin %

                           =  1,18,500 / 34%

                           =  $3,48,529


Required sales = Breakeven point sales - Total revenue

                         = $3,48,529 - 2,54,100

                         = $94,429

                             

answered by: pr balajir rao